OpenStudy (anonymous):

Amanda, Bryce, and Corey enter a race in which they have to run, swim, and cycle over a marked course. Their average speeds are given in the table. Corey finishes first with a total time of 1 hours 45minutes, Amanda comes in second with a time of 2 hours 30 minutes, and Bryce finishes last with a time of 3 hours. Find the distance (in miles) for each part of the race.(hint: use the motion formula d=rt) average speed(mi/h) amanda: 10 running, 4 swimming, and 20 cycling Bryce: 7.5 runnnig, 6swimming,and 15cycling corey: 15 running, 3 swimming,40 cycling a. write the system of equations b.set

OpenStudy (anonymous):

Let d1=distance for running, d2=distance for swimming d3=distance for cycling corey => 15d1+3d2+40d3=1.75 amanda => 10d1+4d2+20d3=2.5 bryce => 7.5d1+6d2+15d3=3

OpenStudy (anonymous):

how would i set this up as a matrix? I know the answers are 5miles running 2 miles swimming and 30 miles cycling. but i dont know how to get that.